5. The money multiplier

In the following data table, calculate the amount of excess reserves and the monetary base and then use the table to answer the questions that follow.

Category

Value

(Billions of Dollars)

Currency, C 450.00
Deposits, D 400.00
Reserves, R 49.00
Required Reserves, RR 46.00
Excess Reserves
Monetary Base

Use the previous data table to fill in the missing cells in the following table. (Hint: Round to the nearest hundredth.)

Ratio

Value

Currency Ratio, kk
Required Reserves Ratio, rrrr   
Excess Reserves Ratio, rexrex
Money Supply Multiplier, MsMMsM   

Based on your answers to the previous questions, the money supply is$
Answer #1

Answer : Table showing information as follows :

Category value in dollars ( Billions)
Currency, C 450.00
Deposits,D 400.00
Reserve ,R 49.00
Required reserve , RR 46.00
Excess reserves= Reserve - Required reserve 3.00
Monetary Base =C+D 850.0

  ​​​​​​ Table showing information about different ratio:

Ratio Value
Currency ratio(kk)=C/D 450/400= 1.125
Required reserve ratio(RR)=RR/D 46/400=0.115
Excess reserve ratio =(ER /D) 3/400=0.0075
Money supply multiplier= 1/r 1/0.115= 8.70

Money supply = Monetary base * Money multipler = 850*8.70= $7395